OpenStudy (anonymous):

If tanθ = 1/3, then secθ = _____.

geerky42 (geerky42):

Knowing that \(\tan\theta = \dfrac{Opp}{Adj}\), it always helps to sketch, like this:

geerky42 (geerky42):

Since \(\sec\theta = \dfrac{Hyp}{Adj}\), you would need to find Hyp using pythagorean theorem

geerky42 (geerky42):

does this help? @garcia22and

OpenStudy (anonymous):

yes! thank you so much!

geerky42 (geerky42):

welcome